# The Geometry of LLM Quantization: GPTQ as Babai’s Nearest Plane Algorithm

### v2

@@ Abstract
Quantizing the weights of large language models (LLMs) from 16-bit to lower bitwidth is the de facto approach to deploy massive transformers onto more affordable accelerators. While GPTQ emerged as one of the standard methods for one-shot post-training quantization at LLM scale, its inner workings are described as a sequence of ad-hoc algebraic updates that obscure geometric meaning or worst-case guarantees. In this work, we show that, when executed back-to-front (from the last to first dimension) for a linear layer, GPTQ is mathematically identical to Babai’s nearest plane algorithm for the classical closest vector problem (CVP) on a lattice defined by the Hessian matrix of the layer’s inputs. This equivalence is based on a sophisticated mathematical argument, and has two analytical consequences: first, the GPTQ error propagation step gains an intuitive geometric interpretation; second, GPTQ inherits the error upper bound of Babai’s algorithm under the assumption that no weights are clipped. Leveraging this bound, we design post-training quantization methods that avoid clipping, and outperform the original GPTQ. In addition, we provide efficient GPU inference kernels for the resulting representation. Taken together, these results place GPTQ on a firm theoretical footing and open the door to importing decades of progress in lattice algorithms towards the design of future quantization algorithms for billion-parameter models.
@@ 1 Introduction
Generative pre-trained transformers (GPT) models contain hundreds of billions of parameters and have massive computational and memory costs (Luccioni et al., 2024). Post-training quantization (PTQ) has emerged as a practical solution for reducing their footprint (Gholami et al., 2021). Among a growing family of methods, GPTQ (Frantar et al., 2023) was the first to push one-shot quantization down to the 4-bit regime, while retaining near-baseline accuracies. GPTQ is still very popular nowadays and yields state-of-the-art results in some regimes (Kurtic et al., 2024).
@@
Our contribution. This paper is the first [^1] The concurrent work of Birnick (2025) appeared on arXiv later than our preprint. to provide a geometric interpretation for GPTQ, which implies a layer-wise global error bound. Our main theoretical results (Section˜4) are (i) the GPTQ optimization problem, i.e. linear-layer quantization with the L2 objective on the output, is equivalent to the closest vector problem (CVP) w.r.t. L2 distance; (ii) the GPTQ algorithm executed from the last to first dimension is the same as Babai’s nearest plane algorithm on the basis of the factorized Hessian matrix, without LLL basis reduction, and this finding holds independently of whether large weights are clipped to the quantization grid (a process known as weight clipping); and (iii) the worst-case layer-wise error in the no-clipping setting is bound tightly by the trace of the diagonal matrix of the LDL decomposition of the Hessian matrix. In addition (Section˜5), we tie our theoretical findings to practical quantization by introducing new no-clipping methods of better accuracy than the original GPTQ, together with efficient GPU inference kernels for the resulting representation.
@@ 2 Related Work
Second-order compression (pruning and quantization). The idea of using Hessian information to guide parameter removal dates back to Optimal Brain Damage (LeCun et al., 1989) and Optimal Brain Surgeon (OBS) (Hassibi et al., 1993). Optimal Brain Compression (OBC) (Frantar & Alistarh, 2022) generalizes OBS to the post-training setting and unifies structured pruning and quantization (also called Optimal Brain Quantizer, OBQ) under a single exact solver. GPTQ (Frantar et al., 2023) inherits OBQ’s error propagation method but applies it in a fixed order, so that the inverse Hessian can be shared and only needs to be computed once. GPTQ only has cubic computational complexity in the column/row dimension, making it suitable for LLMs. QuIP (Chee et al., 2023) proves an error guarantee for GPTQ and proposes the LDLQ method as an equivalent variant of GPTQ.
Lattices, CVP algorithms, and hardness. The closest vector problem (CVP) is NP-complete to approximate within any constant factor under polynomial-time reductions (van Emde Boas, 1981; Micciancio & Goldwasser, 2002; Dinur et al., 2003), motivating decades of approximation algorithms. Babai’s nearest plane heuristic (Babai, 1986) delivers a solution in polynomial time and, when preceded by LLL basis reduction (Lenstra et al., 1982), enjoys a $2^{O\left(n\right)}$ approximation. BKZ basis reduction (Kannan, 1987) further tightens the constant in an exponential-time solver.
@@ 3 Preliminaries and Notations
We use Python-style indexing inside square brackets to select elements and sub-matrices from a tensor, e.g., $[j,:]$ selects the $j$-th row vector, $[:,j]$ selects the $j$-th column vector, and $[j:,j]$ selects the sub-column consisting of rows after $j$-th (included) row in $j$-th column, $[:,J]$ selects the column vectors indexed by set $J$ as a sub-matrix, etc [^2] For more details, please see (NumPy): https://numpy.org/doc/stable/user/basics.indexing.html. The pseudocode for these algorithms is below.
1
Input: $\bm{W},\bm{S},\bm{X},\bm{P},\lambda,\mathbb{Z}_{\dagger}$
Output: $\bm{Z},\bm{Q}$
2
[…]
7for $j\leftarrow 1\ \mathrm{to}\ c$ do
8
9 $\bm{\zeta}\leftarrow\bm{W}[j,:]/\bm{S}[j,:]$
11 $\bm{Q}[j,:]\leftarrow\bm{Z}[j,:]*\bm{S}[j,:]$
12 $\bm{\varepsilon}\leftarrow\bm{Q}[j,:]-\bm{W}[j,:]$
13 $\bm{W}[j:,:]\leftarrow\bm{W}[j:,:]+\bm{L}[j:,j]\bm{\varepsilon}$
14 end for
15
[…]
Algorithm 1 GPTQ
1
Input: $\bm{B},\bm{y}$
Output: $\bm{z}$
2
$\bm{T}\leftarrow\textsc{LLL}\left(\bm{B}\right)$ // transformation
3
$\bm{A}\leftarrow\bm{B}\bm{T}$ // basis reduction
4
$\bm{\Phi}\leftarrow\textsc{QR}\left(\bm{A}\right)$ // orthogonalize
5
[…]
7for $j\leftarrow c\ \mathrm{to}\ 1$ do
8
9 $\zeta\leftarrow\left<\bm{\Phi}[:,j],\bm{y}^{\prime}\right>/\left<\bm{\Phi}[:,j],\bm{A}[:,j]\right$>
10 $\bm{z}[j]\leftarrow\textsc{Round}\left(\zeta,\mathbb{Z}\right)$
11 $\bm{y}^{\prime}\leftarrow\bm{y}^{\prime}-\bm{A}[:,j]\bm{z}[j]$
12 end for
13
[…]
Algorithm 2 Babai’s Nearest Plane
@@ 3.1 Linear-Layer Quantization Problem
Problem. Let $\bm{X}=\left[\bm{x}_{1},\dots,\bm{x}_{n}\right]^{\top}\in\mathbb{R}^{n\times c}$ be the sampled calibration input data of batch size $n$ and input dimension $c$ with $\bm{x}_{i}\in\mathbb{R}^{c}$ and $n\geq c=\mathrm{rank}\left(\bm{X}\right)$. Let $\bm{W}=\left[\bm{w}_{1},\dots,\bm{w}_{r}\right]\in\mathbb{R}^{c\times r}$ be the linear layer weights of input dimension $c$ and output dimension $r$ with $\bm{w}_{i}\in\mathbb{R}^{c}$. Let $\bm{S}=\left[\bm{s}_{1},\dots,\bm{s}_{r}\right]\in\mathbb{R}_{\neq 0}^{c\times r}$ be the non-zero quantization scales with $\bm{s}_{i}\in\mathbb{R}_{\neq 0}^{c}$. Here we consider a general case that applies to any grouping pattern: each weight element $\bm{w}_{i}[j]$ has its own scaling factor $\bm{s}_{i}[j]$. Assume $\bm{S}$ is statically computed using methods like AbsMax or MSE before any weight updates. Let $\mathbb{Z}_{\dagger}\subseteq\mathbb{Z}$ be the quantization grid (representable integers). In the clipping setting, e.g., for INT4 format, $\mathbb{Z}_{\dagger}=\left\{-8,\dots,-1,0,1,\dots,7\right\}$. In the no-clipping setting, $\mathbb{Z}_{\dagger}=\mathbb{Z}$, which allows any integer as the quantization results. Let $\bm{Z}=\left[\bm{z}_{1},\dots,\bm{z}_{r}\right]\in{\mathbb{Z}_{\dagger}}^{c\times r}$ be the (unknown) quantized integers with $\bm{z}_{i}\in\mathbb{Z}_{\dagger}^{c}$. Denote $\bm{Q}=\left[\bm{q}_{1},\dots,\bm{q}_{r}\right]\in\mathbb{R}^{c\times r}$ as the dequantized weights with $\bm{q}_{i}=\mathrm{diag}\left({\bm{s}_{i}}\right)\bm{z}_{i}\in\mathbb{R}^{c}$. The goal is to minimize the L2 error on the layer output $\bm{X}\bm{W}\in\mathbb{R}^{n\times r}$: $\left\|\bm{X}\bm{Q}-\bm{X}\bm{W}\right\|_{2}^{2}=\sum_{i=1}^{r}\left\|\bm{X}\mathrm{\ diag}\left({\bm{s}_{i}}\right)\bm{z}_{i}-\bm{X}\bm{w}_{i}\right\|^{2}$ , i.e, finding $\mathrm{argmin}_{\bm{z}_{i}\in\mathbb{Z}_{\dagger}^{c}}\left\|\bm{X}\mathrm{\ diag}\left({\bm{s}_{i}}\right)\bm{z}_{i}-\bm{X}\bm{w}_{i}\right\|^{2}$ for all $1\leq i\leq r$.
@@
Algorithm˜1 is the pseudocode of GPTQ. The algorithm is identical to the original GPTQ paper (Frantar et al., 2023) except for missing the blocking mechanism that only affects the memory access pattern and computational speed, but not the numerical results. Additional notations are as follows. $\bm{P}\in\left\{0,1\right\}^{c\times c}$ is a permutation matrix that modifies the dimensional order of GPTQ quantization. The default order is front-to-back (from the first to last dimension), i.e., $\bm{P}=\mathbf{I}$. $\lambda\in\mathbb{R}_{+}$ is a small damping factor for computing the Hessian matrix, ensuring the matrix is of full rank. A typical choice is $\lambda=\frac{1}{100c}\sum_{j=1}^{c}\left(\bm{X}^{\top}\bm{X}\right)[j,j]=\frac{1}{100c}\left\|\bm{X}\right\|_{2}^{2}$. Function LDL returns the lower triangular matrix in LDL decomposition. Symbols * and // denote the element-wise multiplication and division.
@@ 3.2 The Closest Vector Problem (CVP)
Problem. Let $\bm{B}=\left[\bm{b}_{1},\dots,\bm{b}_{c}\right]\in\mathbb{R}^{n\times c}$ be a set of $c$ basis vectors of dimension $n$ with $\bm{b}_{j}\in\mathbb{R}^{n}$ and $n\geq c=\mathrm{rank}\left(\bm{B}\right)$. Let $\bm{y}\in\mathbb{R}^{n}$ be an external target vector to approximate. Let $\bm{z}\in\mathbb{Z}^{c}$ be the (unknown) integer vector representing the basis combinations of the lattice vector. The goal is to find the vector on the lattice defined by the basis $\bm{B}$ that is the closest to the target vector $\bm{y}$, i.e., finding $\mathrm{argmin}_{\bm{z}\in\mathbb{Z}^{c}}\left\|\bm{B}\bm{z}-\bm{y}\right\|^{2}$. A visualization of a two-dimensional CVP is shown in Figure˜1 (a).
@@
Figure 1: Upper row: (a) CVP in a two-dimensional lattice; (b) Basis reduction can find a shorter, more orthogonal basis that can potentially improve the results; (c-d) The projection steps in Babai’s nearest plane algorithm. Lower row: rounding boundaries of (e) optimal rounding or Voronoi cells; (f) round-to-nearest (RTN); (g) Babai’s nearest plane algorithm without basis reduction; (h) Babai’s algorithm without basis reduction under the reversely ordered basis.
@@
Babai’s nearest plane algorithm. Babai’s algorithm iteratively projects the target vector onto the nearest hyperplane of a LLL-reduced lattice and rounds the corresponding coefficient. Figure˜1 (b) visualizes the basis reduction step and Figure˜1 (c-d) visualize the projection steps.
@@
Algorithm˜2 is the pseudocode of Babai’s nearest plane algorithm to solve CVP. For better computational efficiency, the pseudocode uses a conceptually equivalent approach. Instead of projecting the target vector to the nearest hyperplane, it moves the target vector along the basis direction towards the hyperplane where the origin lies. The projection error is kept in the updated target vector since it is orthogonal to the hyperplane and will not affect the following projections. Additional notations are as follows. Function LLL returns the transformation matrix of the LLL reduction with parameter delta defaulting to $\frac{3}{4}$. Function QR returns the orthogonal matrix in QR decomposition, the same as the normalized Gram-Schmidt orthogonalization process. $\left$$\cdot$$\cdot\right$> denotes the vector dot product. Function Round is defined as in the GPTQ algorithm.
@@
Babai’s error bound. Figure˜1 shows the rounding boundaries of the optimal (e), round-to-nearest (RTN) (f), and Babai’s algorithm without basis reduction (g-h). Compared to RTN, Babai’s algorithm generates rectangular partitions and thus has a smaller worst-case error. The error bound has been proven in Babai (1986). Formally, let $\bm{\Phi}=\left[\bm{\phi}_{1},\dots,\bm{\phi}_{c}\right]$ be the set of normalized Gram-Schmidt vectors of the LLL-reduced basis $\bm{A}=\left[\bm{a}_{1},\dots,\bm{a}_{c}\right]$. Let $\tilde{\bm{A}}=\left[\tilde{\bm{a}}_{1},\dots,\tilde{\bm{a}}_{c}\right]$ denote the unnormalized Gram-Schmidt vectors with $\tilde{\bm{a}}_{j}=\left<\bm{\phi}_{j},\bm{a}_{j}\right>\bm{\phi}_{j}$. At iteration $j$, the algorithm replaces the exact coefficient $\zeta$ by the closest integer, so the deviation satisfies $\left|\zeta-\bm{z}[j]\right|\leq\frac{1}{2}$. Hence the error component along $\tilde{\bm{a}}_{j}$ has norm at most $\frac{1}{2}\left\|\tilde{\bm{a}}_{j}\right\|$. Because the $\tilde{\bm{A}}$ is orthogonal, these error components add in Euclidean norm, giving a bound on the residual (error) vector $\bm{y}^{\prime}$: $\left\|\bm{y}^{\prime}\right\|^{2}\leq\frac{1}{4}\sum_{j=1}^{c}\|\tilde{\bm{a}}_{j}\|^{2}=\frac{1}{4}\sum_{j=1}^{c}\left<\bm{\phi}_{j},\bm{a}_{j}\right>^{2}$. Babai’s algorithm guarantees to return the center vector of the hyper-cuboid (Figure˜1 (g)) constructed by the unnormalized Gram-Schmidt vectors $\tilde{\bm{A}}$ where the target $\bm{y}$ is located. Equality is attained when the target $\bm{y}$ lies at the corner of the hyper-cuboid, so the bound is tight. Babai (1986) additionally proved a relative error bound for $\gamma$ with $\left\|\bm{B}\bm{z}-\bm{y}\right\|\leq\gamma\cdot\min_{\bm{z}^{\prime}\in\mathbb{Z}^{c}}\left\|\bm{B}\bm{z}^{\prime}-\bm{y}\right\|$. The bound is $1\leq\gamma\leq\sqrt{1+\max_{1\leq j\leq c}\frac{\sum_{j^{\prime}=1}^{j}\left\|\tilde{\bm{a}}_{j^{\prime}}\right\|^{2}}{\left\|\tilde{\bm{a}}_{j}\right\|^{2}}}\leq\sqrt{c+1}\cdot\max_{1\leq j^{\prime}\leq j\leq c}\frac{\left\|\tilde{\bm{a}}_{j^{\prime}}\right\|}{\left\|\tilde{\bm{a}}_{j}\right\|}$.

### v4

@@ Abstract
Quantizing the weights of large language models (LLMs) from 16-bit to lower bitwidth is the de facto approach to deploy massive transformers onto more affordable accelerators. While GPTQ emerged as one of the standard methods for one-shot post-training quantization at LLM scale, its inner workings are described as a sequence of algebraic updates that obscure geometric meaning or worst-case guarantees. In this work, we show that, when executed back-to-front (from the last to first dimension) for a linear layer, GPTQ is mathematically identical to Babai’s nearest plane algorithm for the classical closest vector problem (CVP) on a lattice defined by the Hessian matrix of the layer’s inputs. This equivalence is based on a sophisticated mathematical argument, and has two analytical consequences: first, the GPTQ error propagation step gains an intuitive geometric interpretation; second, GPTQ inherits the error upper bound of Babai’s algorithm under the assumption that no weights are clipped. Leveraging this bound, we design post-training quantization methods that avoid clipping, and outperform the original GPTQ. In addition, we provide efficient GPU inference kernels for the resulting representation. Taken together, these results place GPTQ on a firm theoretical footing and open the door to importing decades of progress in lattice algorithms towards the design of future quantization algorithms for billion-parameter models. Source code is available at https://github.com/IST-DASLab/GPTQ-Babai.
@@ 1 Introduction
Generative pre-trained transformers (GPT) models contain hundreds of billions of parameters and have massive computational and memory costs (Luccioni et al., 2024). Post-training quantization (PTQ) has emerged as a practical solution for reducing their footprint (Gholami et al., 2021). Among a growing family of methods, GPTQ (Frantar et al., 2023) was the first to push one-shot quantization down to the 4-bit regime, while retaining near-baseline accuracies. GPTQ is still very popular nowadays and yields state-of-the-art results in some regimes (Kurtic et al., 2025).
@@
Our contribution. This paper is the first [^1] The concurrent work of Birnick (2025) appeared on arXiv slightly later than our preprint (Chen et al., 2025). to provide a geometric interpretation for GPTQ, which implies a layer-wise global error bound. Our main theoretical results (Section˜4) are (i) the GPTQ optimization problem, i.e., linear-layer quantization with the L2 objective on the output, is equivalent to the closest vector problem (CVP) w.r.t. L2 distance; (ii) the GPTQ algorithm executed from the last to first dimension is the same as Babai’s nearest plane algorithm on the basis of the factorized Hessian matrix, without LLL basis reduction, and this finding holds independently of whether large weights are clipped to the quantization grid (a process known as weight clipping); and (iii) the worst-case layer-wise error in the no-clipping setting is bound tightly by the trace of the diagonal matrix of the LDL decomposition of the Hessian matrix. In addition (Section˜5), we tie our theoretical findings to practical quantization by introducing new no-clipping methods of better accuracy than the original GPTQ, together with efficient GPU inference kernels for the resulting representation.
@@ 2 Related Work
Second-order compression (pruning and quantization). The idea of using Hessian information to guide parameter removal dates back to Optimal Brain Damage (LeCun et al., 1989) and Optimal Brain Surgeon (OBS) (Hassibi et al., 1993). Optimal Brain Compression (OBC) (Frantar and Alistarh, 2022) generalizes OBS to the post-training setting and unifies structured pruning and quantization (also called Optimal Brain Quantizer, OBQ) under a single exact solver. GPTQ (Frantar et al., 2023) inherits OBQ’s error propagation method but applies it in a fixed order, so that the inverse Hessian can be shared and only needs to be computed once. GPTQ only has cubic computational complexity in the column/row dimension, making it suitable for LLMs. QuIP (Chee et al., 2023) proves an error guarantee for GPTQ and proposes the LDLQ method as an equivalent variant of GPTQ.
Lattices, CVP algorithms, and hardness. The closest vector problem (CVP) is NP-complete to approximate within any constant factor under polynomial-time reductions (van Emde Boas, 1981; Micciancio and Goldwasser, 2002; Dinur et al., 2003), motivating decades of approximation algorithms. Babai’s nearest plane heuristic (Babai, 1986) delivers a solution in polynomial time and, when preceded by LLL basis reduction (Lenstra et al., 1982), enjoys a $2^{O\left(n\right)}$ approximation. BKZ basis reduction (Kannan, 1987) further tightens the constant in an exponential-time solver.
@@ 3 Preliminaries and Notations
We use Python-style indexing inside square brackets to select elements and sub-matrices from a tensor, e.g., $[j,:]$ selects the $j$-th row vector, $[:,j]$ selects the $j$-th column vector, and $[j:,j]$ selects the sub-column consisting of rows after the $j$-th (inclusive) row in the $j$-th column, $[:,J]$ selects the column vectors indexed by set $J$ as a sub-matrix, etc [^2] For more details, please see (NumPy): https://numpy.org/doc/stable/user/basics.indexing.html.
@@ 3.1 Linear-Layer Quantization Problem
Problem. Let $\bm{X}=\left[\bm{x}_{1},\dots,\bm{x}_{n}\right]^{\top}\in\mathbb{R}^{n\times c}$ be the sampled calibration input data of batch size $n$ and input dimension $c$ with $\bm{x}_{i}\in\mathbb{R}^{c}$ and $n\geq c=\mathrm{rank}\left(\bm{X}\right)$. Let $\bm{W}=\left[\bm{w}_{1},\dots,\bm{w}_{r}\right]\in\mathbb{R}^{c\times r}$ be the linear layer weights of input dimension $c$ and output dimension $r$ with $\bm{w}_{i}\in\mathbb{R}^{c}$. Let $\bm{S}=\left[\bm{s}_{1},\dots,\bm{s}_{r}\right]\in\mathbb{R}_{\neq 0}^{c\times r}$ be the non-zero quantization scales with $\bm{s}_{i}\in\mathbb{R}_{\neq 0}^{c}$. Here we consider a general case that applies to any grouping pattern: each weight element $\bm{w}_{i}[j]$ has its own scaling factor $\bm{s}_{i}[j]$. Assume $\bm{S}$ is statically computed using methods like AbsMax or MSE before any weight updates. Let $\mathbb{Z}_{\dagger}\subseteq\mathbb{Z}$ be the quantization grid (representable integers). In the clipping setting, e.g., for INT4 format, $\mathbb{Z}_{\dagger}=\left\{-8,\dots,-1,0,1,\dots,7\right\}$. In the no-clipping setting, $\mathbb{Z}_{\dagger}=\mathbb{Z}$, which allows any integer as the quantization results. Let $\bm{Z}=\left[\bm{z}_{1},\dots,\bm{z}_{r}\right]\in{\mathbb{Z}_{\dagger}}^{c\times r}$ be the (unknown) quantized integers with $\bm{z}_{i}\in\mathbb{Z}_{\dagger}^{c}$. Denote $\bm{Q}=\left[\bm{q}_{1},\dots,\bm{q}_{r}\right]\in\mathbb{R}^{c\times r}$ as the dequantized weights with $\bm{q}_{i}=\mathrm{diag}\left({\bm{s}_{i}}\right)\bm{z}_{i}\in\mathbb{R}^{c}$. The goal is to minimize the L2 error on the layer output $\bm{X}\bm{W}\in\mathbb{R}^{n\times r}$: $\left\|\bm{X}\bm{Q}-\bm{X}\bm{W}\right\|_{\mathrm{F}}^{2}=\sum_{i=1}^{r}\left\|\bm{X}\mathrm{\ diag}\left({\bm{s}_{i}}\right)\bm{z}_{i}-\bm{X}\bm{w}_{i}\right\|^{2}$ , i.e, finding $\mathrm{argmin}_{\bm{z}_{i}\in\mathbb{Z}_{\dagger}^{c}}\left\|\bm{X}\mathrm{\ diag}\left({\bm{s}_{i}}\right)\bm{z}_{i}-\bm{X}\bm{w}_{i}\right\|^{2}$ for all $1\leq i\leq r$.
@@
Algorithm˜1 is the pseudocode of GPTQ. The algorithm is identical to the original GPTQ paper (Frantar et al., 2023) except for missing the blocking mechanism that only affects the memory access pattern and computational speed, but not the numerical results. Additional notations are as follows. $\bm{P}\in\left\{0,1\right\}^{c\times c}$ is a permutation matrix that modifies the dimensional order of GPTQ quantization. The default order is front-to-back (from the first to last dimension), i.e., $\bm{P}=\mathbf{I}$. $\lambda\in\mathbb{R}_{+}$ is a small damping factor for computing the Hessian matrix, ensuring the matrix is of full rank. A typical choice is $\lambda=\frac{1}{100c}\sum_{j=1}^{c}\left(\bm{X}^{\top}\bm{X}\right)[j,j]=\frac{1}{100c}\left\|\bm{X}\right\|_{\mathrm{F}}^{2}$. Function LDL returns the lower triangular matrix in the LDL decomposition. Symbols * and // denote element-wise multiplication and division, respectively.
1
Input: original weights $\bm{W}\in\mathbb{R}^{c\times r}$, per-coordinate scales $\bm{S}\in\mathbb{R}^{c\times r}_{\neq 0}$, calibration activation $\bm{X}\in\mathbb{R}^{n\times c}$, permutation $\bm{P}\in\left\{0,1\right\}^{c\times c}$, damping ratio $\lambda>0$, integer grid $\mathbb{Z}_{\dagger}\subseteq\mathbb{Z}$
Output: quantized weights $\bm{Z}\in\mathbb{Z}_{\dagger}^{c\times r}$, dequantized weights $\bm{Q}\in\mathbb{R}^{c\times r}$
2
$\bm{H}\leftarrow\bm{P}^{\top}\left(\bm{X}^{\top}\bm{X}+\lambda\mathbf{I}\right)\bm{P}$ // dampen and reorder Hessian
3
$\bm{L}\leftarrow\textsc{LDL}\left(\bm{H}^{-1}\right)$ // factorize (take the L matrix from the LDL decomposition) the inversed Hessian as the shared coefficients for error propagation
4
$\bm{W},\bm{S}\leftarrow\bm{P}^{-1}\bm{W},\bm{P}^{-1}\bm{S}$ // reorder weights and scales
5
$\bm{Q},\bm{Z}\leftarrow\bm{W},\bm{0}$ // initialize dequantized and quantized weights
6
7for $j\leftarrow 1\ \mathrm{to}\ c$ do
8
 $\bm{\zeta}\leftarrow\bm{W}[j,:]/\bm{S}[j,:]$ // element-wise divide current row by its scales
9
 $\bm{Z}[j,:]\leftarrow\textsc{Round}\left(\bm{\zeta},\mathbb{Z}_{\dagger}\right)$ // quantize coefficients to the target grid
10
 $\bm{Q}[j,:]\leftarrow\bm{Z}[j,:]*\bm{S}[j,:]$ // dequantize current row back to weight space
11
 $\bm{\varepsilon}\leftarrow\bm{Q}[j,:]-\bm{W}[j,:]$ // quantization error for current row
12
 $\bm{W}[j:,:]\leftarrow\bm{W}[j:,:]+\bm{L}[j:,j]\bm{\varepsilon}$ // propagate error to not-yet-quantized rows; broadcast over columns
13
14 end for
15
$\bm{Z},\bm{Q}\leftarrow\bm{P}\bm{Z},\bm{P}\bm{Q}$ // undo reorder to restore original input order; return integers and dequantized weights
16
Algorithm 1 GPTQ
@@ 3.2 The Closest Vector Problem (CVP)
Problem. Let $\bm{B}=\left[\bm{b}_{1},\dots,\bm{b}_{c}\right]\in\mathbb{R}^{n\times c}$ be a set of $c$ basis vectors of dimension $n$ with $\bm{b}_{j}\in\mathbb{R}^{n}$ and $n\geq c=\mathrm{rank}\left(\bm{B}\right)$. Let $\bm{y}\in\mathbb{R}^{n}$ be an external target vector to approximate. Let $\bm{z}\in\mathbb{Z}^{c}$ be the (unknown) integer vector representing the basis combinations of the lattice vector. The goal is to find the vector on the lattice defined by the basis $\bm{B}$ that is closest to the target vector $\bm{y}$, i.e., finding $\mathrm{argmin}_{\bm{z}\in\mathbb{Z}^{c}}\left\|\bm{B}\bm{z}-\bm{y}\right\|^{2}$. A visualization of a two-dimensional CVP is shown in Figure˜1 (a).
@@
Figure 1: Top row: (a) CVP in a two-dimensional lattice; (b) Basis reduction can find a shorter, more orthogonal basis that can potentially improve the results; (c-d) The projection steps in Babai’s nearest plane algorithm. Bottom row: rounding boundaries of (e) optimal rounding or Voronoi cells; (f) round-to-nearest (RTN); (g) Babai’s nearest plane algorithm without basis reduction; (h) Babai’s algorithm without basis reduction under the reversely ordered basis.
@@
Babai’s nearest plane algorithm. Babai’s algorithm iteratively projects the target vector onto the nearest hyperplane of an LLL-reduced lattice and rounds the corresponding coefficient. Figure˜1 (b) visualizes the basis reduction step and Figure˜1 (c-d) visualize the projection steps.
@@
Algorithm˜2 is the pseudocode of Babai’s nearest plane algorithm to solve CVP. For better computational efficiency, the pseudocode uses a conceptually equivalent approach. Instead of projecting the target vector onto the nearest hyperplane, it moves the target vector along the basis direction towards the hyperplane where the origin lies. The projection error is retained in the updated target vector as it is orthogonal to the hyperplane and will not affect subsequent projections. Additional notations are as follows. Function LLL returns the transformation matrix of the LLL reduction with the parameter delta defaulting to $\frac{3}{4}$. Function QR returns the orthogonal matrix in QR decomposition, which is the same as the normalized Gram-Schmidt orthogonalization process. $\left$$\cdot$$\cdot\right$> denotes the vector dot product. Function Round is defined as in the GPTQ algorithm.
1
Input: lattice basis (column vectors) $\bm{B}\in\mathbb{R}^{n\times c}$, target vector $\bm{y}\in\mathbb{R}^{n}$
Output: closest lattice vector’s basis coefficients $\bm{z}\in\mathbb{Z}^{c}$
2
$\bm{T}\leftarrow\textsc{LLL}\left(\bm{B}\right)$ // unimodular transformation matrix from LLL basis reduction
3
$\bm{A}\leftarrow\bm{B}\bm{T}$ // reduce the basis
4
$\bm{\Phi}\leftarrow\textsc{QR}\left(\bm{A}\right)$ // normalized Gram-Schmidt process (take the Q matrix from the QR decomposition)
5
$\bm{y}^{\prime},\bm{z}\leftarrow\bm{y},\bm{0}$ // initialize residual target and integer solution in reduced basis
6
7for $j\leftarrow c\ \mathrm{to}\ 1$ do
8
 $\zeta\leftarrow\left<\bm{\Phi}[:,j],\bm{y}^{\prime}\right>/\left<\bm{\Phi}[:,j],\bm{A}[:,j]\right$> // exact coefficient along the unnormalized Gram-Schmidt vector; ratio between the projections of residual and the reduced basis on the Gram-Schmidt direction
9
 $\bm{z}[j]\leftarrow\textsc{Round}\left(\zeta,\mathbb{Z}\right)$ // round to the nearest plane
10
 $\bm{y}^{\prime}\leftarrow\bm{y}^{\prime}-\bm{A}[:,j]\bm{z}[j]$ // update the residual
11
12 end for
13
$\bm{z}\leftarrow\bm{T}\bm{z}$ // map integer solution back to the original basis and return
14
Algorithm 2 Babai’s Nearest Plane
@@
Babai’s error bound. Figure˜1 shows the rounding boundaries of the optimal (e), round-to-nearest (RTN) (f), and Babai’s algorithm without basis reduction (g-h). Compared to RTN, Babai’s algorithm generates rectangular partitions and thus has a smaller worst-case error. The error bound has been proven in Babai (1986). Formally, let $\bm{\Phi}=\left[\bm{\phi}_{1},\dots,\bm{\phi}_{c}\right]$ be the set of normalized Gram-Schmidt vectors of the LLL-reduced basis $\bm{A}=\left[\bm{a}_{1},\dots,\bm{a}_{c}\right]$. Let $\tilde{\bm{A}}=\left[\tilde{\bm{a}}_{1},\dots,\tilde{\bm{a}}_{c}\right]$ denote the unnormalized Gram-Schmidt vectors with $\tilde{\bm{a}}_{j}=\left<\bm{\phi}_{j},\bm{a}_{j}\right>\bm{\phi}_{j}$. At iteration $j$, the algorithm replaces the exact coefficient $\zeta$ with the closest integer, so the deviation satisfies $\left|\zeta-\bm{z}[j]\right|\leq\frac{1}{2}$. Hence, the error component along $\tilde{\bm{a}}_{j}$ has norm at most $\frac{1}{2}\left\|\tilde{\bm{a}}_{j}\right\|$. Because the $\tilde{\bm{A}}$ is orthogonal, these error components add in Euclidean norm, giving a bound on the residual (error) vector $\bm{y}^{\prime}$: $\left\|\bm{y}^{\prime}\right\|^{2}\leq\frac{1}{4}\sum_{j=1}^{c}\|\tilde{\bm{a}}_{j}\|^{2}=\frac{1}{4}\sum_{j=1}^{c}\left<\bm{\phi}_{j},\bm{a}_{j}\right>^{2}$. Babai’s algorithm guarantees to return the center vector of the hyper-cuboid (Figure˜1 (g)) constructed by the unnormalized Gram-Schmidt vectors $\tilde{\bm{A}}$ where the target $\bm{y}$ is located. Equality is attained when the target $\bm{y}$ lies at the corner of the hyper-cuboid, so the bound is tight. Babai (1986) additionally proved a relative error bound for $\gamma$ with $\left\|\bm{B}\bm{z}-\bm{y}\right\|\leq\gamma\cdot\min_{\bm{z}^{\prime}\in\mathbb{Z}^{c}}\left\|\bm{B}\bm{z}^{\prime}-\bm{y}\right\|$. The bound is $1\leq\gamma\leq\sqrt{1+\max_{1\leq j\leq c}\frac{\sum_{j^{\prime}=1}^{j}\left\|\tilde{\bm{a}}_{j^{\prime}}\right\|^{2}}{\left\|\tilde{\bm{a}}_{j}\right\|^{2}}}\leq\sqrt{c+1}\cdot\max_{1\leq j^{\prime}\leq j\leq c}\frac{\left\|\tilde{\bm{a}}_{j^{\prime}}\right\|}{\left\|\tilde{\bm{a}}_{j}\right\|}$.